Find the General solution of the given system dx/dt=x+2y and dy/dt=4x+3
1
Expert's answer
2020-08-03T18:48:46-0400

We have the following system: "\begin{cases} 
\frac{dx}{dt}=x+2y
\\
\frac{dy}{dt}=4x+3y
\end{cases}" , or in the matrix form: "X^{\prime}=AX," where "X=\begin{bmatrix} x
\\ y
\end{bmatrix}, \ A=\begin{bmatrix} 1 & 2 \\ 4 & 3
\end{bmatrix}" .

The characteristic equation: "\det (A-\lambda I)=\begin{vmatrix} 1-\lambda & 2 \\ 4 & 3-\lambda 
\end{vmatrix}=(1-\lambda )(3-\lambda )-4\times 2=\lambda ^2-4\lambda -5=0."

Eigenvalues: "\lambda _1=-1,\ \lambda _2=5."


Let "v_1" be eigen vector, associated with the eigenvalue "\lambda_1" :

"\begin{bmatrix} 1 -\lambda_1& 2 \\ 4 & 3-\lambda_1
\end{bmatrix} v_1=\begin{bmatrix} 2 & 2 \\ 4 & 4
\end{bmatrix}v_1= \begin{bmatrix} 2v_{11} +2 v_{12}\\ 4 v_{11}+4v_{12}
\end{bmatrix}=0," then "v_1= \begin{bmatrix} 1 \\ -1
\end{bmatrix}" .


Let "v_2" be eigen vector, associated with the eigenvalue "\lambda_2" :

"\begin{bmatrix} 1 -\lambda_2& 2 \\ 4 & 3-\lambda_2
\end{bmatrix} v_2=\begin{bmatrix} -4 & 2 \\ 4 & -2
\end{bmatrix}v_2= \begin{bmatrix} -4v_{21} +2 v_{22}\\ 4 v_{21}-2v_{22}
\end{bmatrix}=0," then "v_2= \begin{bmatrix} 1 \\ 2
\end{bmatrix}" .


Then solution has form "X=c_1e^{\lambda_1 t} v_1+ c_2e^{\lambda_2 t} v_2 = \begin{bmatrix} c_1 e^{-t} +c_2 e^{5t}\\ -c_1e^{-t}+2c_2e^{5t}
\end{bmatrix}" .

Answer: "\begin{cases} x=c_1 e^{-t} +c_2 e^{5t}\\ y=-c_1e^{-t}+2c_2e^{5t}
\end{cases}" .